(15m+3)/(2)=-6
3/2x-15=-12
-7x+3=-15/2
4x+1/2(12x+2)=3
3/2(x+8)=6x

Answer by josgarithmetic(39617)

Examening your work for any mistakes:

(15x+3)/(2)=-6
3/2x-15=-12-------------the first mistake,
-7x+3=-15/2
4x+1/2(12x+2)=3
3/2(x+8)=6x

What you could do:
MULTIPLY BOTH sides by 2;
Add the additive inverse of 3 to both sides;
Multiply the multiplicative inverse of 15 by both sides.
Simplify.
This is your answer.

You basically want to undo everything that has been done to x in order to have prepared the original equation. I very obviously did not carry out the steps. I only stated what steps to perform.